# Chaos and spatial prethermalization in driven-dissipative bosonic chains

### v2

@@ Abstract
Thermalization in quantum many-body systems, the process by which they naturally evolve toward thermal equilibrium, typically unfolds over timescales set by the underlying relaxation mechanisms. Yet, the spatial aspect of thermalization in these systems is less understood. We investigate this phenomenon within the nonequilibrium steady state (NESS) of a Bose-Hubbard chain subject at its boundaries to coherent driving and dissipation, a setup inspired by current designs in circuit quantum electrodynamics. We uncover a two-stage thermalization process along the spatial dimension. Close to the coherent drive, the U(1) symmetry of the phase of the photonic field is restored over a short length scale, while its amplitude relaxes over a much larger scale. This opens up an extensive region of the chain where the photon density remains high, and the chaotic dynamics give rise to a hydrodynamic regime, characterized by local equilibria with a large and slowly-varying effective chemical potential. Dynamical fingerprints of chaos in this NESS are probed using semiclassical out-of-time-order correlators (OTOCs) within the truncated Wigner approximation (TWA). We explore the conditions underlying this protracted thermalization in space and argue that similar prethermal chaotic phases are likely to occur in a broad range of extended driven-dissipative systems.
@@ I Introduction
Thermalization underpins statistical mechanics: generic isolated interacting systems evolve towards thermal equilibrium regardless of their initial conditions. Classically, this is addressed by Boltzmann’s H-theorem which relies on the assumption of molecular chaos [1, 2]. Quantum mechanically, signatures of chaos manifest themselves in the structure of the eigenvalue spectrum and the Eigenstate Thermalization Hypothesis provides a theoretical framework to explain how closed Hamiltonian systems can achieve thermalization under unitary dynamics [3, 4, 5, 6].
Typically, thermalization occurs in two stages. First, through a rapid relaxation of non-conserved quantities to local equilibrium values. Then, the hydrodynamic modes—long-wavelength excitations associated with conserved quantities in the bulk of the system such as energy or particle density—relax by a slower, often diffusive, process. The Fermi-Pasta-Ulam-Tsingou paradox—the apparent lack of thermalization in a nonlinear chain of classical oscillators—posed an unexpected challenge to this understanding  [7]. It is now widely accepted that this system’s proximity to an integrable point is responsible for prolonged transient dynamics [8]. In quantum mechanics, long-lived transients are similarly observed in near-integrable systems. When an integrable system, characterized by many conserved quantities, undergoes a quench in which an integrability-breaking pertubation is suddenly turned on, it initially relaxes to a quasi-equilibrium state described by a Generalized Gibbs Ensemble [9, 10, 11, 12, 13]. The subsequent relaxation to a true thermal state can be exceedingly slow, especially when the integrability-breaking interaction is weak [14]. Such slow thermalization dynamics are also seen in periodically-driven systems, where high-frequency driving lead to the emergence of long-lived “prethermal” states [15, 16, 17, 18, 19]. These states retain memory of the initial conditions through approximately conserved quantities, and theoretical proposals and experiments have been put forward to leverage these time-dependent protocols for quantum state engineering [20, 21, 22, 23, 24].
Figure 1: Schematic of the boundary-driven dissipative Bose-Hubbard chain and key results. (a) Tight-binding array of $L$ nonlinear resonators described by the Hamiltonian in Eq. (2) subject to a drive of amplitude $F$ coherently injecting photons at the leftmost site and to single-photon losses at both ends of the chain. The interplay of interaction, drive, and dissipation leads to a nonequilibrium steady state (NESS). (b) Lyapunov growth between neighboring Wigner trajectories are used to identify chaotic dynamics. (c) In the chaotic regime, the chain hosts three distinct domains illustrated by their steady-state local Wigner functions $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$: a nonsymmetric nonthermal domain near the left boundary, an extensive prethermal phase where the U(1)-symmetry of the phase is restored and which hosts a high density of photons only saturated by the Kerr non-linearity, and a U(1)-symmetric thermal domain near the right boundary which is characterized by thermal fluctuations over the vacuum state.
With recent advances in quantum device fabrication, these theoretical considerations have now gained significant experimental relevance for quantum technologies and can now be addressed by specifically engineering many-body systems. In particular, in circuit quantum electrodynamics, nonlinear oscillators are routinely constructed by coupling microwave resonators to nonlinear elements such as Josephson junctions. These building blocks can be assembled into long chains of nonlinear oscillators with up to nearly 100 sites, as demonstrated in studies such as Refs. [25, 26, 27], and they can be effectively described by Bose-Hubbard models. Similar chains have been realized in other quantum architectures, including trapped ions [28], semiconductor micropillars [29], and ultracold gases in optical lattices [30].
These devices are inherently prone to intrinsic losses as well as extrinsic dissipation channels introduced by measurement apparatuses. Furthermore, they are operated away from thermal equilibrium by processes such as gate sequences in quantum computers, pulsed signals in quantum optimal control, or continuous wave drives in quantum optical setups. This open, nonequilibrium nature poses significant conceptual challenges: established concepts such as chaos, thermalization, and the potential for persistent prethermal states must be reconsidered in this context of driven-dissipative dynamics. Pioneering works with classical systems, such as the boundary-driven dissipative Klein-Gordon chain, have demonstrated that the interplay between drive, dissipation, and interactions can create a complex dynamical landscape, including unconventional transport phenomena [31, 32, 33], which are simply absent in their isolated counterparts [34].
@@
In this work, we investigate the route to thermalization in quantum chains of nonlinear bosonic modes. Photons are coherently injected at one end of the chain, and the coupling to the input feedline and output channels brings incoherent photon losses at both ends of the chain. We leverage the fact that these systems operate in regimes well described by semiclassical approaches to promote well-established concepts and tools from classical chaos theory to analyze the ergodic properties of the nonequilibrium steady state (NESS) of the open quantum chain. Specifically, we use the truncated Wigner approximation (TWA) and out-of-time-order correlators (OTOCs) to follow the transition from regular to chaotic regimes. In the latter, we examine the thermalization along the chain, spatially resolving the transition from the strongly nonequilibrium state near the coherent drive to the expected thermal state at the opposite end of the chain. Between these extremes, we identify an extended prethermal phase where the U(1) symmetry, initially broken by the boundary drive, is restored and the chain locally equilibrates to high-density states characterized by a large and slowly-varying effective chemical potential. We attribute this phenomenon to a significant mismatch between the short relaxation scale of the phase degree of freedom and the longer hydrodynamic relaxation of the amplitude sector of the photonic field. Our findings are directly relevant to current experimental platforms, and we propose diagnostics based on routinely measured quantities, which can be determined through quantum-state tomography via, e.g., heterodyne detection.
@@ II.1 Boundary-driven dissipative Bose-Hubbard chain
Here, $\hat{a}_{\ell}^{\dagger}$ ($\hat{a}_{\ell}$) are the bosonic creation (annihilation) operators for the photons in the $\ell$-th resonator mode of frequency $\omega_{0}$. $\Delta:=\omega_{\rm d}-\omega_{0}$ is the pump-to-resonator detuning, $J$ is the hopping amplitude between neighboring resonators, and $U$ is the strength of the onsite Kerr nonlinearity. Our model is motivated by state-of-the-art experimental cavity or circuit QED setups [25, 26, 27] with a weak but finite interaction, $|U|\ll|\Delta|,|J|$. While $U=0$ makes the model trivially integrable, a finite $U$ ensures the nonintegrability of the Bose-Hubbard Hamiltonian on sizeable chains [35]. $F$ is the amplitude of the driving field, which explicitly breaks the U(1)-symmetry of the Bose-Hubbard chain. We take $U>0$, $\Delta>0$, $J>0$, and $F\geq 0$ ($U<0$, $\Delta<0$, $J<0$, and $F\leq 0$ yields identical results).
The Lindblad dissipators at sites $\ell=1$ and $\ell=L$ are defined as
@@
where $\hat{L}_{\ell}=\sqrt{\gamma}\hat{a}_{\ell}$ are the local jump operators modeling incoherent one-photon losses to the environment at a rate $\gamma>0$. When nondriven and isolated, $F=\gamma=0$, the Bose-Hubbard chain is far from the Mott insulating regime, and it is naturally expected to thermalize (see, e.g.[36, 37]). In the presence of drive and dissipation, it is expected to reach a unique NESS, $\hat{\rho}^{\rm ss}:=\lim\limits_{t\to\infty}\hat{\rho}(t)$, carrying a uniform DC current of photons from the left drive to the right drain. The initial conditions to Eq. (1) are inconsequential to our results and, unless stated otherwise, we simply use the vacuum state $\hat{\rho}(0)=\bigotimes_{\ell=1}^{L}|0\rangle_{\ell}\langle 0|_{\ell}$ where $|0\rangle_{\ell}$ is the Fock state with no excitation in the $\ell$-th resonator. We take $\gamma$ as the unit of energy and set $\Delta=2.5$, $J=2$, and $U=0.1$ for the remainder of this study.
@@
${\rm i}\frac{\partial\alpha_{\ell}}{\partial t}$ $=-f(\alpha_{\ell})-J(\alpha_{\ell-1}+\alpha_{\ell+1})\,,\ l=2,...,L-1$
@@
Although the TWA is exact only for quadratic models ($U=0$), it has been successfully applied to describe dissipative phase transitions, disordered systems, time crystals, and quantum chaos in a variety of weakly nonlinear driven-dissipative systems [40, 41, 42, 43]. We justify the use of the TWA in our model by the weakness of the Kerr nonlinearity—a condition easily achievable in current circuit QED architectures [44]; under this condition, the TWA faithfully describe the NESS of single or multiple nonlinear driven resonators [40] and boundary-driven dimers [43]. We further validate this approach by benchmarking it against exact results for small system sizes and comparing it with other approximation schemes (see the Supplementary Information). The choice of the TWA is motivated by the considerable reduction of the computational complexity it offers: the exponential growth of the Hilbert space in Eq. (1) is cut down to a linear growth in the number of the stochastic equations (II.1), thus enabling direct numerical integration of long chains of resonators up to times past the transient dynamics and deep in the steady-state regime. Furthermore, Wigner trajectories extend the classical notion of phase-space trajectories to the quantum regime, thereby offering a practical means to investigate the exponential sensitivity to initial conditions (or the lack thereof), which is a constitutive hallmark of chaotic dynamics [45]. Notably, while the density matrix and the local Wigner functions remain constant in the NESS, individual Wigner trajectories keep fluctuating and can thus be used to analyze the integrable versus chaotic character of the dynamics in the NESS [46, 47].
Figure 2: Steady-state quantum chaos via phase OTOC. Dynamics of the steady-state phase OTOC $D_{1,\ell}^{\rm ss}(\tau)$, defined below Eq. (5), computed at sites $\ell=1,10,20,...,80$ of an $L=80$ chain and for two representative drive strengths $F$. (a) $F=7.5$: Exponential growth at a Lyapunov rate $\lambda\approx 2.8$ and ballistic spreading at a butterfly velocity $v=2J$, signaling chaotic dynamics. (b) $F=12$: Sub-exponential growth, signaling non-chaotic or regular dynamics. Dashed lines indicate the chaotic bound on the saturation value of the phase OTOC. The results are computed upon averaging over $5\times 10^{3}$ independent Wigner trajectories. Throughout the manuscript, the dissipation rate $\gamma$ is taken as the unit of energy unit and the other parameters are set to $\Delta=2.5$, $J=2$, and $U=0.1$.
II.2 Regular and chaotic regimes in the NESS
Dynamical signatures of quantum chaos are commonly probed by means of out-of-time-order correlation functions such as average square commutators of the form $\textrm{Tr}\left([\hat{W}_{\ell}(\tau),\hat{V}_{k}]^{\dagger}[\hat{W}_{\ell}(\tau),\hat{V}_{k}]\hat{\rho}\right)$ where $\hat{V}_{k}$ and $\hat{W}_{\ell}$ are operators with local support around sites $k$ and $\ell$, respectively [48, 49, 50]. We focus on the OTOC between the number and the phase degrees of freedom of the resonators, i.e., $\hat{n}_{k}$ and $\hat{\phi}_{\ell}$ evaluated at time $t$ and later time $t+\tau$, respectively. In the Methods, we show that the semiclassical formulation of this phase OTOC can be computed as [51, 52, 53, 54, 55, 56, 57]
$D_{k,\ell}(t,\tau):=1-\left\langle\cos\left[\phi_{\ell}^{(a)}(t+\tau)-\phi_{\ell}^{(b)}(t+\tau)\right]\right\rangle.$ (5)
Here, $\phi_{\ell}(t):=\textrm{arg}\,\alpha_{\ell}(t)$ is the phase of the complex field in the $\ell$-th resonator along an individual Wigner trajectory. The superscripts $(a)$ and $(b)$ refer to two replicas of the system which are identical until time $t$ at which an infinitesimal local perturbation is applied at site $k$ in replica $b$: $\alpha_{k}^{(b)}(t)\to\alpha_{k}^{(b)}(t)+\varepsilon$. The subsequent evolution is computed using the same quantum noise realization for the two replicas. delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over realizations of the noise and the limit $\varepsilon\to 0$ is implied.
Figure 3: Nonequilibrium steady-state phase diagram. Steady-state properties of the last resonator in the chain, $\ell=L$, as a function of the chain length $L$ and drive strength $F$. (a) Saturation value of the steady-state phase OTOC, $D^{\rm ss}_{1,L}(\tau\to\infty)$, defined below Eq. (5). (b) Photon number $n^{\rm ss}_{L}$. (c) Photon-number fluctuation $\delta n_{L}$ defined in Eq. (6). (d-f) Normalized local Wigner functions, $\tilde{W}^{\rm ss}_{L}(\alpha,\alpha^{*}):=W_{L}^{\rm ss}(\alpha,\alpha^{*})/\textrm{max}[W^{\rm ss}_{L}(\alpha,\alpha^{*})]$, for chain lengths $L=10$ (red colorscale) and $L=80$ (blue half-width contours) and for drive strengths (d) $F=2$, (e) $F=7.5$, and (f) $F=12$, corresponding to the markers in panel (a). The three distinct regimes labelled by (I), (II) and (III) are discussed in the text. Results are computed upon averaging over $10^{3}$ independent Wigner trajectories. Statistics are further improved by averaging over a time window $\Delta\tau$ after the NESS is reached, with $\Delta\tau=25$ for $D^{\rm ss}_{1,L}(\tau\to\infty)$, and $\Delta\tau=500$ for $n^{\rm ss}_{L}$ and $\delta n^{\rm ss}_{L}$. The other parameters are set as in Fig. 2.
We compute the steady-state phase OTOC, defined as $D^{\rm ss}_{k,\ell}(\tau):=\lim\limits_{t\to\infty}D_{k,\ell}(t,\tau)$, to map the nonequilibrium steady-state phase diagram of the boundary-driven dissipative Bose-Hubbard chain. This quantity successfully captures basics features of quantum information spreading in extended many-body systems [58] and we show in the Methods that $D^{\rm ss}_{1,\ell}(\tau)$ clearly exhibits a causal light-cone structure with a ballistic spreading of information characterized by a butterfly velocity $v=2J$ [59]. On general grounds, we have $D_{k,\ell}^{\rm ss}(\tau=0)=0$, $D_{k,\ell}^{\rm ss}(\tau>0)$ is expected to increase with $\tau$ whenever trajectories in the two replicas start deviating, and, eventually, to saturate to a finite value $D_{k,\ell}^{\rm ss}(\tau\to\infty)$. Both the growth and the saturation regimes of $D_{k,\ell}^{\rm ss}(\tau)$ shed light on the chaotic versus regular nature of the dynamics in the NESS. In chaotic dynamics, the typical distance between two trajectories with nearly identical initial conditions is expected to grow exponentially – a hallmark of Lyapunov instability, which is often taken a defining feature of classical chaos. Quantum mechanically, this exponential growth is expected to be captured by OTOCs in systems that have a well-defined semiclassical limit or large-$N$ limit. The saturation value $D_{k,\ell}^{\rm ss}(\tau\to\infty)$ is related to the square of the available local phase space volume. $D_{k,\ell}^{\rm ss}(\tau\to\infty)\approx 0$ corresponds to situations where the phases in replica $a$ and $b$ remain strongly correlated, suggestive of regular dynamics. In contrast, in chaotic regimes where the trajectories decorrelate rapidly, $\phi^{(a)}$ and $\phi^{(b)}$ can be seen as uniformly distributed random phases, yielding a saturation value at its chaotic bound, $D^{\rm ss}_{k,\ell}(\tau\to\infty)\approx 1$.
In Fig. 2, we illustrate the growth of the steady-state phase OTOC for two representative values of the drive strength $F$. In panel (a), $D^{\rm ss}_{1,\ell}(\tau)$ shows a rapid exponential growth of the form $D^{\rm ss}_{1,\ell}(\tau)\sim\exp[\lambda(t-l/v)]$ where $\lambda$ is a Lyapunov rate, followed at scrambling times on the order of $1/2J$ by a saturation regime where $D^{\rm ss}_{1,\ell}(\tau\to\infty)\approx 1$, i.e., maximal decorrelation. Altogether, the semiclassical phase OTOC $D_{1,\,\ell}^{\rm ss}(\tau)$ captures both the Lyapunov growth and the saturation regimes expected of quantum chaotic dynamics. In panel (b), $D^{\rm ss}_{1,\ell}(\tau)$ instead displays early-time oscillations and the overall growth is slower than exponential. Eventually, the late-time saturation value $D^{\rm ss}_{1,\ell}(\tau)$ is significantly less than $1$. This is indicative of regular dynamics.
Beyond the two representative values of $F$ used in Fig. 2, we verified across the examined parameter range that $D^{\rm ss}_{1,\ell}(\tau\to\infty)=1$ consistently corresponds to exponential growth, while $D^{\rm ss}_{1,\ell}(\tau\to\infty)\ll 1$ corresponds to sub-exponential growth. Therefore, we use the saturation value of the steady-state phase OTOC, $D^{\rm ss}_{1,\ell}(\tau\to\infty)$, as a proxy to map the chaotic versus regular character of the nonequilibrium steady-state dynamics as the drive strength $F$ and the chain length $L$ are varied. Additionally, we monitor two standard quantities in quantum optics, namely the steady-state photon number $n_{\ell}^{\rm ss}:=\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}$ and its fluctuations $\Delta n_{\ell}^{\rm ss}:=\langle(\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell})^{2}\rangle_{\textrm{ss}}-\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\textrm{ss}}^{2}$ by means of the quantity
$\delta n^{\rm ss}_{\ell}:=\Delta n^{\rm ss}_{\ell}-n^{\rm ss}_{\ell}$ (6)

### v3

# Chaotic and quantum dynamics in driven-dissipative bosonic chains

@@ Abstract
Thermalization in quantum many-body systems typically unfolds over timescales governed by intrinsic relaxation mechanisms. Yet, its spatial aspect is less understood. We investigate this phenomenon in the nonequilibrium steady state (NESS) of a Bose-Hubbard chain subject to coherent driving and dissipation at its boundaries, a setup inspired by current designs in circuit quantum electrodynamics. The dynamical fingerprints of chaos in this NESS are probed using semiclassical out-of-time-order correlators (OTOCs) within the truncated Wigner approximation (TWA). At intermediate drive strengths, we uncover a two-stage thermalization along the spatial dimension: phase coherence is rapidly lost near the drive, while amplitude relaxation occurs over much longer distances. This separation of scales gives rise to an extended hydrodynamic regime exhibiting anomalous temperature profiles, which we designate as a “prethermal” domain. At stronger drives, the system enters a nonthermal, non-chaotic finite-momentum condensate characterized by sub-Poissonian photon statistics and a spatially modulated phase profile, whose stability is undermined by quantum fluctuations. We explore the conditions underlying this protracted thermalization in space and argue that similar mechanisms are likely to emerge in a broad class of extended driven-dissipative systems.
@@ I Introduction
Chaos and thermalization are ubiquitous features in classical and quantum many-body systems. Classical chaotic dynamics are well understood in terms of Lyapunov instability: two initially nearby trajectories in phase space diverge exponentially in time [1]. In quantum mechanics, however, the notion of phase-space trajectory is lost, and the characterization of chaos becomes more subtle. Quantum chaotic dynamics are typically diagnosed through the statistical properties of the energy spectrum [2] as well as through the temporal behavior of out-of-time-ordered correlation functions [3]. With the recent advances of quantum computing platforms, especially within circuit quantum electrodynamics (QED), the study of chaos in quantum devices has gained significant relevance [4, 5, 6, 7, 8, 9]. State-of-the art driven-dissipative circuit QED architectures are nowadays able to control tens of coupled nonlinear oscillators, enabling the experimental realization of atom-photon bound states [10, 11, 12], two-dimensional Bose-Hubbard lattices [13], quantum chaotic and thermalized models [14, 15, 16] and concatenated bosonic qubits [17]. In the ongoing effort to fabricate and operate increasingly complex systems—featuring a growing number of individual degrees of freedom and enhanced nonlinearities, chaotic dynamics and thermalization are both an opportunity for quantum simulations and a threat for the coherent manipulation of quantum information.
Large, isolated and nonintegrable many-body systems are generically expected to thermalize at sufficiently long times. Classically, the microscopic understanding of thermalization is addressed by Boltzmann’s H-theorem, which relies on the assumption of molecular chaos [18, 19]. Quantum mechanically, the eigenstate thermalization hypothesis provides a theoretical framework to explain how closed Hamiltonian systems can achieve thermalization under unitary dynamics [20, 21, 22, 2, 23]. Typically, thermalization occurs in two stages. First, non-conserved quantities rapidly relax to local equilibrium values. Then, the hydrodynamic modes—long-wavelength excitations associated with conserved quantities such as energy or particle density—relax through a slower, often diffusive, process. This prolonged transient before complete thermalization is particularly pronounced in the proximity of integrability, where it has sometimes been dubbed prethermal regime. This is notably the case when an integrability-breaking perturbation is turned on suddenly [24, 25, 26] or in a periodic fashion [27, 28, 29, 30, 31]. The system initially relaxes to a quasi-equilibrium state [32, 33], and subsequently converges to a true thermal state on an exceedingly long timescale [34].
Quantum devices are inherently prone to intrinsic losses as well as extrinsic dissipation channels introduced by measurement, and their open, nonequilibrium nature poses significant conceptual challenges. While dissipative quantum chaos has gained considerable attention [35, 36, 37, 38, 39, 40, 41, 42], thermalization in chaotic dissipative systems remains only partially explored. In this work, we investigate the route to thermalization in driven-dissipative chains of nonlinear bosonic modes, where photons are coherently injected at one end of the chain, and the coupling to the input and output channels induces incoherent photon losses at both ends of the chain. This is a prototypical and versatile model for quantum hardware based on superconducting circuits, within reach of current experimental setups [43, 44, 45]. Similar chains have been realized in other quantum architectures, including trapped ions [46], semiconductor micropillars [47], and ultracold gases in optical lattices [48].
At the classical level, i.e., neglecting the quantum fluctuations, recent studies of boundary-driven dissipative Klein-Gordon chains have revealed a rich phenomenology, including unconventional transport behaviors [49, 50, 51], which are absent in their isolated counterparts [52]. As the drive strength is increased, three distinct nonequilibrium steady-state (NESS) regimes have been identified: First, a quasi-linear regime in which nonlinearities are essentially irrelevant, the dynamics remain regular, and transport is ballistic. Second, a hydrodynamic regime characterized by local thermal equilibria and superdiffusive transport. At even stronger drives, this regime gives way to a resonant nonlinear wave (RNW) regime, characterized by a spatially periodic phase profile and ballistic transport. The RNW regime has been shown to remain stable under weak but finite thermal fluctuations induced by the environments at the two boundaries of the chain.
@@
In this article, we explore the structure of the quantum many-body NESS and we investigate how quantum fluctuations impact this rich dynamical landscape. We leverage the fact that these systems are experimentally operated in regimes well described by semiclassical approaches to promote well-established concepts and tools from classical chaos theory to analyze the ergodic properties of the nonequilibrium steady state (NESS) of the open quantum chain. Specifically, we use the truncated Wigner approximation (TWA) and out-of-time-order correlators (OTOCs) to follow the transition from regular to chaotic regimes beyond the classical descriptions of Refs. [49, 50, 51]. In the chaotic regime, we spatially resolve the transition from the strongly nonequilibrium state near the coherent drive to the low-temperature state expected at the opposite end of the chain. Between these extremes, we identify an extended phase, referred to as the prethermal phase, in which the $\mathbb{U}(1)$ symmetry, initially broken by the boundary drive, is restored, and the chain locally equilibrates to high-density states. Notably, this phase exhibits anomalous heating: the temperature increases with the distance from the boundary where energy is injected, resulting in a temperature gradient that opposes both the photon and energy currents. We attribute this phenomenon to a significant mismatch between the short relaxation scale of the phase degree of freedom and the longer hydrodynamic relaxation of the amplitude sector of the photonic field. Away from the hydrodynamic regime, we find that quantum effects significantly impact the RNW regime. In particular, quantum fluctuations demote the long-range coherence of the phase modulation and can even destabilize the RNW regime in sufficiently long chains. Our findings are directly relevant to current experimental platforms, and we propose diagnostics based on routinely measured quantities, which can be determined through quantum-state tomography via, e.g., heterodyne detection.
The paper is structured as follows. In Sec. II.1, we introduce our model of boundary driven-dissipative bosonic chains and the TWA approach. In Sec. II.2, we describe the steady-state phase diagram as a function of the strength of the boundary drive. In Sec. II.3, we detail the hydrodynamic regime and its prethermal phase. In Sec. II.4, we analyze the RNW regime along with its distinct quantum features. We conclude in Sec. III by outlining potential directions for future work.
Figure 1: Boundary-driven dissipative Bose-Hubbard chain: schematics and key results. (a) Tight-binding array of $L$ nonlinear resonators described by the Hamiltonian in Eq. (2) subject to a drive of amplitude $F$ coherently injecting photons at the leftmost site and to single-photon losses at both ends of the chain. The interplay of interaction, drive, and dissipation leads to a nonequilibrium steady state (NESS). (b) Lyapunov growth between neighboring Wigner trajectories is used to identify chaotic dynamics. (c) In the chaotic regime, the chain hosts three distinct domains illustrated by their steady-state local Wigner functions $W_{\ell}(\alpha,\alpha^{*})$: a nonsymmetric nonthermal domain near the left boundary, an extensive prethermal phase where the $\mathbb{U}(1)$ symmetry of the phase is restored and which hosts a high density of photons only saturated by the Kerr non-linearity, and a $\mathbb{U}(1)$-symmetric thermal domain near the right boundary which is characterized by fluctuations over the vacuum state.
@@ II.1 Boundary-driven dissipative Bose-Hubbard chain
Here, $\hat{a}_{\ell}^{\dagger}$ ($\hat{a}_{\ell}$) are the bosonic creation (annihilation) operators for the photons in the $\ell$-th resonator mode of frequency $\omega_{0}$. $\Delta:=\omega_{\rm d}-\omega_{0}$ is the pump-to-resonator detuning, $J$ is the hopping amplitude between neighboring resonators, and $U$ is the strength of the onsite Kerr nonlinearity. Our model is motivated by state-of-the-art experimental cavity or circuit QED setups [43, 44, 45] with a weak but finite interaction, $|U|\ll|\Delta|,|J|$. While $U=0$ makes the model trivially integrable, a finite $U$ ensures the non-integrability of the Bose-Hubbard Hamiltonian on sizable chains [53]. Unlike in the classical Klein-Gordon chains studied in Refs. [50, 51], the Kerr nonlinearity features a $\mathbb{U}(1)$ symmetry associated with photon number conservation in the bulk of the chain. $F$ is the amplitude of the driving field, which explicitly breaks this symmetry at one end of the chain. We take $U>0$, $\Delta>0$, $J>0$, and $F\geq 0$ ($U<0$, $\Delta<0$, $J<0$, and $F\leq 0$ yields identical results). The Lindblad dissipators at sites $\ell=1$ and $\ell=L$ are defined as
@@
where $\hat{L}_{\ell}=\sqrt{\gamma}\hat{a}_{\ell}$ are the local jump operators modeling incoherent one-photon losses to the zero-temperature environment at a rate $\gamma>0$. When not driven and isolated, i.e., $F=\gamma=0$, the Bose-Hubbard chain is far from the Mott insulating regime, and it is naturally expected to thermalize (see, e.g., Refs. [54, 55]). In the presence of drive and dissipation, it is expected to reach a unique NESS, $\lim\limits_{t\to\infty}\hat{\rho}(t)$, carrying uniform DC currents of photons and energy flowing from the left drive to the right drain. As initial conditions for Eq. (1), we simply choose the vacuum state, $\hat{\rho}(0)=\bigotimes_{\ell=1}^{L}|0\rangle_{\ell}\langle 0|_{\ell}$, where $|0\rangle_{\ell}$ denotes the Fock state with zero excitation in the $\ell$-th resonator. We take $\gamma$ as the unit of energy and set $\Delta=2.5$, $J=2$, and $U=0.1$ for the remainder of this study.
@@
${\rm i}\frac{\partial\alpha_{\ell}}{\partial t}$ $=-f(\alpha_{\ell})-J(\alpha_{\ell-1}+\alpha_{\ell+1})\,,\ \ell=2,...,L-1$
@@
Although the TWA is exact only for quadratic models ($U=0$), it has been successfully applied to describe dissipative phase transitions, disordered systems, time crystals, and quantum chaos in a variety of weakly nonlinear driven-dissipative systems [63, 64, 65, 66, 67, 68, 69, 70]. We justify the use of the TWA in our model by the weakness of the Kerr nonlinearity—a condition easily achievable in current circuit QED architectures [71]; under this condition, the TWA faithfully describes the NESS of single or multiple nonlinear driven resonators [66] and boundary-driven dimers [69]. We further validate this approach by benchmarking it against exact results for small system sizes and comparing it with other approximation schemes (see the Supplementary Information). The choice of the TWA is motivated by the considerable reduction of the computational complexity it offers: the exponential growth of the Hilbert space in Eq. (1) is reduced to a linear growth in the number of stochastic equations (II.1), thus enabling direct numerical integration of long chains of resonators up to times past the transient dynamics and deep in the steady-state regime. Furthermore, Wigner trajectories extend the classical notion of phase-space trajectories to the quantum regime, thereby offering a practical means of investigating the exponential sensitivity to initial conditions (or lack thereof), which is a constitutive hallmark of chaotic dynamics [1]. Notably, while the density matrix and local Wigner functions remain constant in the NESS, individual Wigner trajectories keep fluctuating and can thus be used to analyze the integrable versus chaotic character of the dynamics in the NESS [38, 42].
II.2 Dynamical regimes in the NESS
We begin by investigating the NESS phase diagram as a function of the drive amplitude $F$ and the chain length $L$. Unless specified otherwise, all expectation values reported below are computed in the NESS. To characterize the different steady-state regimes, we extract observables at the last site of the chain. We monitor two standard quantities in quantum optics that can easily be probed within the TWA framework: the steady-state photon number $n_{\ell}:=\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle$ and the variance of its fluctuations $(\Delta n_{\ell})^{2}:=\langle(\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell})^{2}\rangle-\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle^{2}$ by means of the quantity
$\delta n_{\ell}:=(\Delta n_{\ell})^{2}-n_{\ell}$ (5)